# Model Predictive Inferential Control of Neural State-Space Models for Autonomous Vehicle Motion Planning

### v3

@@ VII-B The Overtaking Scenario
MPIC-X w. $10$ particles $18,992$ $0.029$ $17.31$  &  $3.71$ $-86.26$  &  $-99.59$
MPIC-X w. $50$ particles $18,923$ $0.065$ $13.65$  &  $0.48$ $-69.10$  &  $-99.12$
MPIC-X w. $80$ particles $18,778$ $0.091$ $12.77$  &  $-0.29$ $-56.82$  &  $-98.72$
@@
MPIC-X w. $10$ particles $18,064$ $0.055$ $-7.36$  &  $4.97$ $-83.4$  &  $-99.75$
MPIC-X w. $50$ particles $17,633$ $0.141$ $-9.57$  &  $2.46$ $-58.04$  &  $-99.37$
MPIC-X w. $80$ particles $17,473$ $0.196$ $-10.39$  &  $1.53$ $-41.74$  &  $-99.13$
@@
Next, we compare the MPIC-X algorithm against gradient-based MPC in terms of computation time and planning performance. To solve gradient-based MPC, we use two solvers: MATLAB’s fmincon with the default interior point method, and CasADi [70] with IPOPT [71].
@@ IX Conclusion
We apply the MPIC-X algorithm to highway driving scenarios via extensive simulations. The simulation results validate the capability of the MPIC-X algorithm in dealing with the control of NSS models for motion planning. Its computation is very fast and almost insensitive to neural network architectures, while well applicable and scalable to long prediction horizons, compared to gradient-based MPC. The proposed framework and algorithm hold a potential for addressing various other robotics and engineering problems that involve control of machine learning models.

### v4

@@ VII-B The Overtaking Scenario
[…]
MPIC-X w. $50$ particles $18,923$ $0.065$ $13.65$ $-0.48$ $-69.10$  &  $-99.12$
MPIC-X w. $80$ particles $18,778$ $0.091$ $12.77$ $-0.29$ $-56.82$  &  $-98.72$
@@
MPIC-X w. $10$ particles $18,064$ $0.055$ $-7.36$ $4.97$ $-83.40$  &  $-99.75$
MPIC-X w. $50$ particles $17,633$ $0.141$ $-9.57$ $2.46$ $-58.04$  &  $-99.37$
MPIC-X w. $80$ particles $17,473$ $0.196$ $-10.4$ $1.53$ $-41.74$  &  $-99.13$
@@
Next, we compare the MPIC-X algorithm against gradient-based MPC in terms of computation time and planning performance. To solve gradient-based MPC, we use two solvers: MATLAB’s fmincon with the default interior point method, and CasADi [74] with interior point optimizer (IPOPT) [75].
@@ IX Conclusion
We apply the MPIC-X algorithm to highway driving scenarios via extensive simulations. The simulation results validate the capability of the MPIC-X algorithm in dealing with the control of NSS models for motion planning. Its computation is very fast and almost insensitive to neural network architectures, while well applicable and scalable to long prediction horizons, compared to gradient-based MPC. The proposed framework and algorithm hold a potential for addressing various other robotics and engineering problems that involve the control of machine learning models.